OpenStudy (anonymous):

I need some help.. I've already done these problems and i got them back and they need fixed.. will someone please help me

OpenStudy (anonymous):

find the surface area of the solids. the pyramids are regular and the prisms, cylinders, and cones are right

OpenStudy (anonymous):

For the first one, you can find the surface area of the cylinder part ( the part that looks like a can) by using 2(pi)r * height

OpenStudy (anonymous):

Then add the area of the bottom circle (pi)r^2

OpenStudy (anonymous):

and then add SA of the cone, pi*r*s s is the slant height

OpenStudy (anonymous):

thats what i did and i got 211.95cmin^2

OpenStudy (anonymous):

ok, let me try and work it out with the numbers, just a sec

OpenStudy (anonymous):

Do you know the right answer?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

I got 203.47cm^2

OpenStudy (anonymous):

wekkm 203.58 cm^2 when I use the pi button

OpenStudy (anonymous):

oops, I mean "well"

OpenStudy (anonymous):

for the cylinder, it's 30.6pi, the circle base is 9pi, and the cone is 25.2pi

OpenStudy (anonymous):

so it would be 30.6pi+9pi+25.2pi

OpenStudy (anonymous):

yes, and with that I got 64.8pi

OpenStudy (anonymous):

For the 2nd picture, there are 6 sides that are rectangles, so that would be 6(base*height). I can't tell in the picture, is the height of each triangle 5, or what does the 5 go to?

OpenStudy (anonymous):

to the slant eight of the triangle

OpenStudy (anonymous):

height*

OpenStudy (anonymous):

6 there are 6 triangles, each with a base of 3 and slant height of 5.....that would be 6 * 1/2 (base * height)

OpenStudy (anonymous):

So, 6(3*6) +6(1/2)(3*5)

OpenStudy (anonymous):

and then would you have to add the area of the base?

OpenStudy (anonymous):

855+64.8pi

OpenStudy (anonymous):

where did you get 855?

OpenStudy (anonymous):

6(3*6) +6(1/2)(3*5)

OpenStudy (anonymous):

6(3*12) +6(1/2)(3*5) that would be a 12 in the first ( )

OpenStudy (anonymous):

153

OpenStudy (anonymous):

64.8pi went to the first drawing...the one above is for the 2nd picture

OpenStudy (anonymous):

I get 261 in^2

OpenStudy (anonymous):

i thought the first pic was 203.58cm^2

OpenStudy (anonymous):

yes, that's what I got....it was 64.8pi, which = 203.58cm^2

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

for the 2nd drawing I get 261in^2 + area of that hexagon base

OpenStudy (anonymous):

261 in^2 so this is the second pic surface area?

OpenStudy (anonymous):

yes, but I dont' know how to find area of the base

OpenStudy (anonymous):

will u help me with some other ones that ive already done i just got them wrong

OpenStudy (anonymous):

i will try

OpenStudy (anonymous):

okay.. for this one i needed to find the value of x.. and i got 3

OpenStudy (anonymous):

v=1/3 pi r^2 h

OpenStudy (anonymous):

I get 3 also

OpenStudy (anonymous):

so redoing it i get \[56.5=1/3\times \pi \times x ^{2}\times6\]

OpenStudy (anonymous):

3 meters

OpenStudy (anonymous):

\[56.5=6.28x ^{2}\]

OpenStudy (anonymous):

yes, i get that too

OpenStudy (anonymous):

Did it count off because you didn't put meters?

OpenStudy (anonymous):

maybe

OpenStudy (anonymous):

oh, wait, is x the diameter? We only found the radius.

OpenStudy (anonymous):

so x = 6 meters

OpenStudy (anonymous):

56.5/6.28=8.996

OpenStudy (anonymous):

yeah, then sq rt of that gives the radius of that circle as 3 meters

OpenStudy (anonymous):

but in the drawing, the x represents the diameter, which would be twice the radius, or 6 meters

OpenStudy (anonymous):

I've got to go....good luck!

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

you're welcome